OpenStudy (anonymous):

Simplify. 6(3x - 5) + 2(3x - 5) - 8(3x - 5) A)0 B)24x - 40 C)36x - 60 D)48x - 80

OpenStudy (anonymous):

use the distributive property

OpenStudy (anonymous):

6(3x - 5) + 2(3x - 5) - 8(3x - 5) = 18x -30 +6x -10 -24x +40 = 18x +6x -24x -30 -10 +40 = 0

OpenStudy (anonymous):

thanks :]]

OpenStudy (anonymous):

@gabriel11 Openstudy.com/code-of-conduct

OpenStudy (anonymous):

explain this equation

OpenStudy (anonymous):

@gabriel11 Easy... Distribute 6(3x - 5) then 2(3x - 5) and -8(3x - 5) after that Combine all like terms.

OpenStudy (anonymous):

okey